WATERWAYS CORPORATION
Budgeted Costs for Irrigation Projects
Time
Charges
Material
Loading
Charges
Labor wages (5,700 hours) $238,260
Supervisor’s salary $62,000
Clerical and accountant wages 58,140 4,000
Irrigation supplies manager 39,000
Overhead 53,010 21,720
   Total $349,410 $126,720


Waterways has budgeted for 5,700 labor hours. It desires a $13 profit margin per hour of labor and 15% profit on materials. It estimates the total invoice cost of materials in the coming year will be $660,000.

A.Compute the rate per hour of labor. (Round answer to 2 decimal places, e.g. 52.75.)
labor rate $74.3

B.Compute the material loading charge.

material loading charge = 35.20% (this is wrong)

C.

Waterways has received a request for a bid to do a parkway for the city. The irrigation manager estimates that it will take about a month to complete the project and require 530 hours of labor and $84,000 of materials. Compute the total estimated bid for the parkway project. (Round answer to 0 decimal places, e.g. 5,275.)

Total estimated bid for the parkway project $ ???????
Answer #1

Ans:

A. Calculate the rate per hour of labor as shown below:

Rate Per Hour of labor

=Total Direct labor cost/Total direct labor hours+Desired Margin per labour Hour

=$349,410/5700+13

=61.3+13

=74.30

Therefore the direct labor rate per hour is $74.30

B. Calculate the Material loading charge as shown below:

Material loading charges=Total material Loading charges/Estimated cost of materials + Desired Margin on Materials

=$126,720/$660,000+15%

=.192+.15

=.342

=34.20%

The Material Loading charge is 34.20%

C. Calculate the estimated bid for the Project

Estimated Bid=Materials costs +Labor costs+ Margin on materials and labor

=84,000+84,000+34.20/530 Hours*74.30

=112,728+39,379

=$152,107

The estimated bid for the Project is $152,107